# Theory of local $\mathbb{Z}_{2}$ topological markers for finite and periodic two-dimensional systems

### v1

@@ VI.1 Choosing the trial projections for the LZ2M
The topological obstruction arising when $C\neq 0$ manifests in the overlap matrix between projected states (the $S$ matrix of Eq. 15), that becomes singular somewhere in the BZ [5], prohibiting the construction of MLWFs. As discussed in Sec. III.2, in order to have well-defined local $\mathbb{Z}_{2}$ topological invariants based on TR symmetry, we split the occupied manifold in two TR-conjugate sub-manifolds and select trial projection orbitals |$\ket{g_{n}}$ that break TR symmetry and minimize the spillage $\gamma$ between $\mathcal{P}_{\Theta}=\mathcal{P}_{1}+\Theta\mathcal{P}_{1}\Theta^{-1}$ and the ground-state projector $\mathcal{P}$. If we were to choose trial projections localized on just one site of the cell, the information on the other basis site would be missing since TR symmetry acts only on spin in real space, resulting in a large value of $\gamma$. Nonetheless, the resulting WFs can still have contributions from all the lattice sites, even though the trial projections are localized on different sites [52]. To minimize the spillage, we observe that it is convenient to choose initial projections with contributions from all sites. In the Kane-Mele model, this can be realized, for instance, by these two choices:
@@
where |$\ket{\mathbf{R},A,\sigma}$ is a normalized $\delta$-like orbital with spin $\sigma$ centered on site $A$ in the primitive cell defined by $\mathbf{R}$. Here we use the lattice vectors $\mathbf{R}$ of the pristine system defined in the primitive cell, as opposed to the lattice vectors of the supercell $\mathbf{\tilde{R}}$. Indeed, the supercell on which the boundary conditions are imposed is constructed first through repetition of its pristine primitive cell. Then, we can choose the trial projections just in the primitive cell and replicate the choice across the whole supercell, even in the disordered case. The spillage between the projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}$ can then be measured as a function of the linear size of the system $L$. The results we obtain are reported in Fig. 1, where we compare the spillage 11footnotemark: 1 of the projections defined in Eqs. 29 and 30 with the choice made in Ref. [52], that is
@@
The spillage between the projections defined in Eq. 31 is small compared to other choices, we but since the trial functions are localized on a single site, the marker we can compute with those functions will not work. The projections of Eq. 30 are localized on both sites but are characterized by a large spillage, so we can expect a poor convergence of the marker with this choice. Overall, the best choice seems to be the one of Eq. 29 and will be used from now on when referring to the LZ2Ms: these orbitals mix all sites resulting in the smallest spillage. When smearing is introduced in the calculation of the LZ2Ms, additional trial projections must be chosen to account for the partly occupied states. This case is analogous to that of a metal, where WFs can be computed via a “disentanglement” procedure [55]. However, it looks unclear how the additional projections should be chosen, and how the procedure should be carried out to ensure the requirements described before. Indeed, one would have to localize the trial functions in an arbitrary unit cell in such a way as to preserve the exponential localization of the projectors and to make the marker independent with respect to the specific cell the trial functions are localized in. Hence, the construction of a LZ2M with smearing is left to further investigation.
@@ VI.2 Convergence
In Figure 2, we show the convergence of the LSCMs and LZ2Ms, evaluated on a unit cell (two sites) in the center of the supercell, as a function of the supercell linear size $L$ in both OBCs and PBCs. For the PBC markers, we report in Fig. 2 the convergence of the symmetric formulations only, as they converge faster than the asymmetric ones. The comparison between the PBC asymmetric and symmetric formulas in reported Fig. 3, that shows clearly the faster convergence of the symmetric formulations. The convergence of the PBC markers is polynomial in both the trivial and the topological phases since the error is dominated by the approximation of the derivative with finite differences in the single-point limit [33]. Within OBCs, we can see that the convergence of the LSCM is always exponential (due to the exponential localization of the projectors $P_{\pm}$ [37]), while the LZ2M converges exponentially in the trivial phase but polynomially in the topological one. This different behavior is due to the fact that, in the topological phase, the projector computed by first obtaining qWFs and then taking half of them is exponentially localized only at short distances, while its tail at large distances decays slower. This can be clearly seen in Fig. 4, that shows, for the topological phase, the decay of the matrix elements of the projectors used in the OBC LSCM and OBC LZ2M in the bulk of a $100\times 30$ crystallite with 6000 sites. We could not determine unambiguously whether this long-tail is exponential or polynomial, that would require studying much larger systems. In the trivial phase, the projectors we obtain are always exponentially localized, resulting in an exponential convergence of the markers. We note that the LZ2M converges to the expected value despite its projectors not being fully exponentially localized. Indeed, being the invariant a topological and a ground-state quantity, we only need the information about the local electronic structure to retrieve its local value in real space. For this reason, we argue that the projectors need to be exponentially localized only at short distances, while the tail behavior determines other aspects such as the convergence of the formula with the system size. We attribute the slower decay of the tails of $\mathcal{P}_{1}$ to be a side effect of minimizing the spillage $\gamma$ [Eq. 18] between the projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}$. Because of the topological obstruction, we cannot obtain TR symmetric and exponentially localized WFs. In our procedure, we require that the sum of the projector computed with half qWFs and its TR partner is as close as possible to the initial projector, still performing a minimal breaking of TR symmetry. That, in turn, results in an almost TR symmetric gauge that leads to qWFs that cannot be truly exponentially localized because of the topological obstruction. In other words, we are looking for a compromise between the breaking of TR symmetry and the exponential localization of WFs that allows us to probe the local $\mathbb{Z}_{2}$ topology in real space. As a result, the spread [Eq. 16] of qWFs displays a very slow divergence with the system size, and the long-tail behavior of $\mathcal{P}_{1}$ could be attributed to the fact that the qWFs we obtain are not truly exponentially localized.
@@ VI.3 Disordered systems
In Figure 5, we show that both the LSCM and LZ2M are capable of charting the local topology also in presence disorder. In particular, we consider Anderson disorder through a random on-site potential uniformly distributed in the interval $[-W/2,W/2]$, where $W$ is the disorder amplitude [56]. For small disorder amplitudes the topology of the system is expected to survive, hence we consider a Kane-Mele model in its topological phase ($\Delta/\lambda_{SO}=0.5$, $\lambda_{R}/\lambda_{SO}=1$) with $W=2$, for which the single-point spin-Chern number [12] predicts a topological phase. As shown in Fig. 5, the PBC markers display the expected local topology aside from small fluctuations due to the disordered environment. To account for the lack of periodicity, a macroscopic average on a radius $R=3$ (in units of the lattice parameter) has been employed. Specifically, the absence of boundaries of the supercell is evident when comparing the PBC markers with the OBC ones. In fact, despite showing the same behavior inside the bulk of the system, near the boundary the PBC markers are continuous while the OBC ones are not, due to the presence of metallic edge states. Within OBCs the individual LCMs ±$\mathfrak{C}_{\pm}(\mathbf{r})$ are such that their trace over the whole sample vanishes, so the edge states contribution to the marker should compensate the non-trivial bulk topology. When computing the LSCM and LZ2M, since they are defined only modulo $2$, the divergence of the individual LCMs results in a discontinuous behavior of the $\mathbb{Z}_{2}$ markers.
@@ VI.4 Heterojunctions and superlattices
Last, we validate and compare the performance of the local markers on trivial/topological heterojunctions and superlattices. We compute the LSCM in both OBCs [Fig. 6] and PBCs [Fig. 7] for a $6000$-site supercell made of alternating topological ($\Delta/\lambda_{SO}=0.5$) and trivial ($\Delta/\lambda_{SO}=8$) regions. Both the LSCM and LZ2M are able to chart the local topology, however, since the LSCM demonstrates a better convergence and performance with respect to the LZ2M we use only the former to discuss the numerical results for inhomogeneous systems. The presence of metallic edge states is highlighted in OBCs both at the edge of the supercell and at the interface between the subsystems, while in PBCs they appear only at the interface, being the superlattice periodic in both directions. In the calculation of the PBC marker, we set the smearing temperature $T_{s}=0.05$ to improve convergence.
@@ VII Summary and Conclusions
We developed a framework to probe the local $\mathbb{Z}_{2}$ topology of 2D TR-symmetric systems in real space for both finite samples in OBCs and extended systems in PBCs, by introducing a number of topological markers. All markers are based on the fundamental idea that the occupied manifold can be split into two TR-conjugate subspaces, where a corresponding individual local Chern marker can then be calculated. In the first approach (LSCM), the separation is performed by diagonalizing the projected spin operator $\mathcal{P}S_{z}\mathcal{P}$, as its spectrum is generally made of two sectors of positive and negative eigenvalues separated by a gap, in the spirit of Prodan’s spin-Chern number [37]. Hence, the LSCM depends on the existence of such spectral gap, which is not an essential property of QSHIs but seems to be usually satisfied also in presence of rather strong Rashba SOC: we never observe such a gap closure for the $\mathcal{P}S_{z}\mathcal{P}$ spectrum in our simulations on the Kane-Mele model, at least as long as the Hamiltonian gap remains finite. Notably, the LSCM could in principle be used also for systems with a small breaking of TR symmetry, as in the TR-symmetry-broken quantum spin Hall effect [57], provided that the aforementioned spectral gap remains finite, since the construction does not explicitly make use of TR.
@@
The LSCM is computationally easier to implement and use, essentially avoiding the need of qWFs or any other localization procedure, so we suggest using that method in applications whenever possible. Otherwise, if the system of interest is such that the projected spin operator does not have a finite gap, the LZ2M approach can always be used (if TR symmetry holds). We emphasize that the construction of the LZ2Ms is more delicate, as an accurate choice of the initial projection orbitals is needed to obtain well-behaved topological markers. The specific projections we selected in our numerical experiments might be improved and the choice is not universal, depending on the details of the system and its topological phase.
@@
Our local markers are based on the ground-state electron distribution only, hence being very suited to large-scale ab initio electronic structure simulations of non-crystalline systems, not only under Anderson disorder but also in presence of defects or interfaces, and for amorphous topological materials [63, 64, 65] or quasi-crystals [66, 14]; in all cases even if TR symmetry or the perpendicular component of the spin, $S_{z}$, are not conserved.
@@ VIII ACKNOWLEDGMENTS
N.B. acknowledges a scholarship from the Department of Physics of the University of Trieste and Collegio Universitario “Luciano Fonda”; A.M. acknowledges support from the ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union - NextGenerationEU – PNRR, Missione 4 Componente 2 Investimento 1.4.

### v2

@@ VI.1 Choosing the trial projections for the LZ2M
The topological obstruction arising when $C\neq 0$ manifests in the overlap matrix between projected states (the $S$ matrix of Eq. 15), that becomes singular somewhere in the BZ [5], prohibiting the construction of MLWFs. As discussed in Sec. III.2, in order to have well-defined local $\mathbb{Z}_{2}$ topological invariants based on TR symmetry, we split the occupied manifold in two TR-conjugate sub-manifolds and select trial projection orbitals |$\ket{g_{n}}$ that break TR symmetry and minimize the spillage $\gamma$ between $\mathcal{P}_{\Theta}=\mathcal{P}_{1}+\Theta\mathcal{P}_{1}\Theta^{-1}$ and the ground-state projector $\mathcal{P}$. If we were to choose trial projections localized on just one site of the cell, the information on the other basis site would be missing since TR symmetry acts only on spin in real space, resulting in a large value of $\gamma$. Nonetheless, the resulting WFs can still have contributions from all the lattice sites, even though the trial projections are localized on different sites [58]. To minimize the spillage, we observe that it is convenient to choose initial projections with contributions from all sites. In the Kane-Mele model, this can be obtained, for instance, with these two choices:
@@
where |$\ket{\mathbf{R},A,\sigma}$ is a normalized $\delta$-like orbital with spin $\sigma$ centered on site $A$ in the primitive cell defined by $\mathbf{R}$. Here we use the lattice vectors $\mathbf{R}$ of the pristine system defined in the primitive cell, as opposed to the lattice vectors of the supercell $\mathbf{\tilde{R}}$. Indeed, the supercell on which the boundary conditions are imposed is constructed through repetitions of its pristine primitive cell. Then, we can choose the trial projections just in the primitive cell and replicate the choice across the whole supercell, even in the disordered case. The spillage between the projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}$ can then be measured as a function of the linear size of the system $L$. The results we obtain are reported in Fig. 1, where we compare the spillage 11footnotemark: 1 of the projections defined in Eqs. 29 and 30 with the choice made in Ref. [6], that is
@@
The spillage between the projections (Pr3) defined in Eq. 31 is small compared to other choices, we but since the trial functions are localized on a single site, the marker we can compute with those functions will not work. The projections (Pr2) [Eq. 30] are localized on both sites but are characterized by a large spillage, so we can expect a poor convergence of the marker with this choice. Overall, the best choice is (Pr1) [Eq. 29] and it will be used from now on when referring to the LZ2Ms: These orbitals mix all sites resulting in the smallest spillage. In general, to identify suitable initial projections, one could first compute the MLWFs and then mix them to minimize the spillage. When smearing is introduced in the calculation of the LZ2Ms, additional trial projections must be chosen to account for the partly occupied states. This case is analogous to that of a metal, where WFs can be computed via a “disentanglement” procedure [61]. However, it looks unclear how the additional projections should be chosen, and how the procedure should be carried out to ensure the requirements described before. Indeed, one would have to localize the trial functions in an arbitrary unit cell in such a way as to preserve the exponential localization of the projectors and to make the marker independent with respect to the specific cell the trial functions are localized in. Hence, the construction of a LZ2M with smearing is left to further investigation.
@@ VI.2 Convergence
In Figure 2, we show the convergence of the LSCMs and LZ2Ms, evaluated on a unit cell (two sites) in the center of the supercell, as a function of the supercell linear size $L$ in both OBCs and PBCs. For the PBC markers, we report in Fig. 2 the convergence of the symmetric formulations only, as they converge faster than the asymmetric ones. The comparison between the PBC asymmetric and symmetric formulas is reported in Fig. 3, that shows clearly the faster convergence of the symmetric formulations. The convergence of the PBC markers is polynomial in both the trivial and the topological phases since the error is dominated by the approximation of the derivative with finite differences in the single-point limit [35]. Within OBCs, we can see that the convergence of the LSCM is always exponential (due to the exponential localization of the projectors $P_{\pm}$ [42]), while the LZ2M converges exponentially in the trivial phase but polynomially in the topological one. This different behavior is due to the fact that, in the topological phase, the projector is exponentially localized only at short distances: Here, the projector is computed by first obtaining qWFs and then taking half of them. This can be clearly seen in Fig. 4, that shows, for the topological phase, the decay of the matrix elements of the projectors used in the OBC LSCM and OBC LZ2M in the bulk of a $100\times 30$ crystallite with 6000 sites. We could not determine unambiguously whether this long tail is exponential or polynomial, that would require studying much larger systems. In the trivial phase, the projectors are always exponentially localized, resulting in an exponential convergence of the markers. We note that the LZ2M converges to the expected value despite its projectors not being fully exponentially localized. Indeed, since the invariant is a topological and a ground-state quantity, we only need the information about the local electronic structure to retrieve its local value in real space. For this reason, we argue that the projectors need to be exponentially localized only at short distances, while the tail behavior determines other aspects such as the convergence of the formula with the system size. We attribute the slower decay of the tails of $\mathcal{P}_{1}$ to be a side effect of minimizing the spillage $\gamma$ [Eq. 18] between the projectors $\mathcal{P}$ and $\mathcal{P}_{\Theta}$. Because of the topological obstruction, we cannot obtain TR symmetric and exponentially localized WFs. In our procedure, we require that the sum of the projector computed with half qWFs and their TR partners is as close as possible to the initial projector, still performing a minimal breaking of TR symmetry. That, in turn, results in an almost TR symmetric gauge that leads to qWFs that cannot be truly exponentially localized because of the topological obstruction. In other words, we are looking for a compromise between the breaking of TR symmetry and the exponential localization of WFs that allows us to probe the local $\mathbb{Z}_{2}$ topology in real space. As a result, the spread [Eq. 16] of qWFs displays a very slow divergence with the system size, and the long-tail behavior of $\mathcal{P}_{1}$ could be attributed to the fact that the qWFs we obtain are not truly exponentially localized.
@@ VI.3 Disordered systems
In Figure 5, we show that both the LSCM and LZ2M are capable of charting the local topology also in presence disorder. In particular, we consider Anderson disorder through a random on-site potential uniformly distributed in the interval $[-W/2,W/2]$, where $W$ is the disorder amplitude [62]. For weak disorder the topology of the system is expected to survive, hence we consider a Kane-Mele model in its topological phase ($\Delta/\lambda_{SO}=0.5$, $\lambda_{R}/\lambda_{SO}=1$) with $W=2$, for which the single-point spin-Chern number [12] predicts a topological phase. As shown in Fig. 5, the PBC markers display the expected local topology aside from small fluctuations due to the disordered environment. To account for the lack of periodicity, a macroscopic average on a radius $R=3$ (in units of the lattice parameter) has been employed. Specifically, the absence of boundaries of the supercell is evident when comparing the PBC markers with the OBC ones. In fact, despite showing the same behavior inside the bulk of the system, near the boundary the PBC markers are continuous while the OBC ones are not, due to the presence of metallic edge states. Within OBCs the individual LCMs ±$\mathfrak{C}_{\pm}(\mathbf{r})$ are such that their trace over the whole sample vanishes, so the edge states contribution to the marker should compensate the non-trivial bulk topology. When computing the LSCM and LZ2M, since they are defined only modulo $2$, the divergence of the individual LCMs results in a discontinuous behavior of the $\mathbb{Z}_{2}$ markers.
@@ VI.4 Heterojunctions and superlattices
Last, we validate and compare the performance of the local markers on trivial/topological heterojunctions and superlattices. We compute the LSCM in both OBCs [Fig. 6] and PBCs [Fig. 7] for a $6000$-site supercell made of alternating topological ($\Delta/\lambda_{SO}=0.5$) and trivial ($\Delta/\lambda_{SO}=8$) regions. Both the LSCM and LZ2M are able to chart the local topology, however, since the LSCM demonstrates a better convergence and performance with respect to the LZ2M we use only the former to discuss the numerical results for inhomogeneous systems. The presence of metallic edge states is highlighted in OBCs both at the edge of the supercell and at the interface between the subsystems, while in PBCs they appear only at the interface, as the superlattice is periodic in both directions. In the calculation of the PBC marker, we set the smearing temperature $T_{s}=0.05$ to improve convergence. Within OBCs, the Bianco-Resta LCM [21] vanishes when the trace is taken over the entire sample, and the topological marker evaluated at the metallic edges (where the Chern number is ill-defined) diverges. For QSHIs, the same holds for the local individual Chern markers ±$\mathfrak{C}_{\pm}(\mathbf{r})$. Here, since the local spin-Chern marker is well-defined only modulo $2$, the divergence of ±$\mathfrak{C}_{\pm}(\mathbf{r})$ results in the discontinuity observed at the interfaces between regions with different topology.
@@ VII Summary and Conclusions
We developed a framework to probe the local $\mathbb{Z}_{2}$ topology of 2D TR-symmetric systems in real space for both finite samples in OBCs and extended systems in PBCs, by introducing a number of topological markers. All markers are based on the fundamental idea that the occupied manifold can be split into two TR-conjugate subspaces, where a corresponding individual local Chern marker can then be calculated. In the first approach (LSCM), the separation is performed by diagonalizing the projected spin operator $\mathcal{P}S_{z}\mathcal{P}$, as its spectrum is generally made of two sectors of positive and negative eigenvalues separated by a gap, in the spirit of Prodan’s spin-Chern number [42]. Hence, the LSCM depends on the existence of such spectral gap, which is not an essential property of QSHIs but seems to be usually satisfied also in presence of rather strong Rashba SOC: We never observe such a gap closure for the $\mathcal{P}S_{z}\mathcal{P}$ spectrum in our simulations on the Kane-Mele model, at least as long as the Hamiltonian gap remains finite. However, as discussed in Ref. [33], 3D strong topological insulators are characterized by a vanishing gap of $\mathcal{P}S\mathcal{P}$ for all choices of the spin operator $S$. Notably, the LSCM could in principle be used also for systems with a small breaking of TR symmetry, as in the TR-symmetry-broken quantum spin Hall effect [63], provided that the aforementioned spectral gap remains finite, since the construction does not explicitly make use of TR.
@@
The LSCM is computationally easier to implement and use, essentially avoiding the need of qWFs or any other localization procedure, so we suggest using that method in applications whenever possible. Otherwise, if the system of interest is such that the projected spin operator does not have a finite gap, the LZ2M approach can always be used (if TR symmetry holds). We emphasize that the construction of the LZ2Ms is more delicate, as an accurate choice of the initial projection orbitals is needed to obtain well-behaved topological markers. The specific projections we selected in our numerical experiments might be improved, and the choice is not universal, depending on the details of the system and its topological phase.
@@
Our local markers are based on the ground-state electron distribution only, hence being very suited to large-scale ab initio electronic structure simulations of non-crystalline systems, not only under Anderson disorder but also in presence of defects or interfaces, and for amorphous topological materials [69, 70, 71] or quasi-crystals [72, 14]; in all cases even if TR symmetry or the perpendicular component of the spin, $S_{z}$, are not conserved. Moreover, the construction of our markers could potentially also be applied to the study of topological phonons [73] and mechanical-optical setups [74], where the dynamical matrix plays a role similar to the electronic Hamiltonian.
@@ VIII ACKNOWLEDGMENTS
N.B. acknowledges a scholarship from the Department of Physics of the University of Trieste and Collegio Universitario “Luciano Fonda”; A.M. acknowledges that this study was partially funded by the European Union - NextGenerationEU, through the ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing – (CUP J93C22000540006, PNRR Investimento M4.C2.1.4). The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.